# QUOTA: Quantifying Objects with Text-to-Image Models for Any Domain

### v1

@@ Abstract
We tackle the problem of quantifying the number of objects by a generative text-to-image model. Rather than retraining such a model for each new image domain of interest, which leads to high computational costs and limited scalability, we are the first to consider this problem from a domain-agnostic perspective. We propose QUOTA, an optimization framework for text-to-image models that enables effective object quantification across unseen domains without retraining. It leverages a dual-loop meta-learning strategy to optimize a domain-invariant prompt. Further, by integrating prompt learning with learnable counting and domain tokens, our method captures stylistic variations and maintains accuracy, even for object classes not encountered during training. For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of object quantification accuracy and adaptability across unseen domains in text-to-image generation. Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new benchmark for efficient and scalable text-to-image generation for any domain.
@@ 1 Introduction
Text-to-image generative models [67, 40, 57, 7] have achieved remarkable success in generating high-quality images from textual descriptions. However, accurately controlling the count of objects in these generated images remains challenging, particularly when faced with text prompts that embody domains different from the ones seen during training. Recently, Zafar et al. [66] introduce prompt learning to refine the object counting prompts. While effective in controlled scenarios, this method struggles to maintain accurate object counts when confronted with varying domains. As a result, achieving domain-generalizable object quantification remains a crucial and unresolved challenge across various generative models, including diffusion and flow-matching models. Figure 1 illustrates this issue, showing how standard generative models like SDXL [41] struggle with accurate object quantification across different domains, whereas our approach provides consistent and accurate results across diverse domains.
Figure 1: Quantifying objects for any domain. The images outlined in red represent cases where the generated images have incorrect object counts. SDXL [41] struggles to control object counts across various domains, demonstrating inconsistencies even in domains typically encountered during training. IoCo by Zafar et al. [66], trained specifically on the photo domain, also faces challenges in generalizing object quantification to unseen styles. Our proposed QUOTA method achieves consistent and accurate object quantification for any domain.
@@
Domain generalization in machine learning aims to create models that perform well on unseen domains, reducing the need for retraining and enabling greater adaptability across diverse settings. A wide range of approaches has been developed for this purpose, including methods that focus on learning domain-invariant features and leveraging diverse training environments [31, 9, 63]. Among these approaches, meta-learning has emerged as a promising technique, enabling models to adapt to new tasks by utilizing prior knowledge from related tasks [20, 60, 12]. Meta-learning has demonstrated success in improving domain generalization across various applications by optimizing models to adapt to different scenarios. Despite these advancements, there has been no exploration of meta-learning for domain-generalizable object quantification in text-to-image generation. In this paper, we fill this gap by proposing a novel meta-learning framework specifically designed to enable robust object quantification across diverse and unseen domains within text-to-image models.
@@
Specifically, we make the following three contributions in this paper. First, we introduce a domain-agnostic setting for object quantification in generative text-to-image models, emphasizing the need for precise, domain-invariant object quantification across domains. This setting is essential for adapting models to diverse and unseen prompts without retraining, addressing both scalability and efficiency challenges. Second, we propose a dual-loop meta-learning strategy for prompt optimization, designed to make prompts domain-invariant. In the training phase, we divide text prompts into meta-train and meta-test domains. In the inner loop, we optimize prompt parameters on the meta-train domain to meet specific domain requirements. The outer loop refines these parameters on the meta-test domain, allowing the model to learn a prompt that generalizes across unseen domains. Third, we incorporate prompt learning with learnable quantification and domain tokens, enabling the model to capture stylistic variations effectively while preserving object quantification accuracy across domains. This approach allows for robust adaptation to new domains and unseen object classes without modifying the model architecture or retraining.
For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of accuracy and adaptability across unseen domains in text-to-image generation.
@@
Comparative evaluations against state-of-the-art methods demonstrate that QUOTA outperforms existing models in both object quantification accuracy and semantic alignment, setting a new benchmark for efficient, adaptable, and scalable text-to-image generation.
@@ 2 Related Work
The field of image generation has seen extensive study, from early advancements with GANs [15, 34, 42, 43, 59, 30, 68] to recent breakthroughs with diffusion models [19, 6, 45, 13, 24]. Diffusion models, in particular, have unlocked remarkable capabilities for generating high-quality, diverse images directly from natural language descriptions, exemplified by models like DALL-E 2 [46], Imagen [51], Parti [65], Make-A-Scene [13], Stable Diffusion (SD) [49], and CogView2 [8]. Our method builds on text-conditioned image generation, specifically exploring how to enhance object counting consistency across domains. Below, we review related areas in editing and controlling image generation models.
Personalized Image Generation. Personalization in image generation typically focuses on fine-tuning models to represent specific concepts from limited examples. Textual Inversion  [14], for instance, learns a new token to represent a concept using just 3-5 images, while DreamBooth [50] fine-tunes diffusion model parameters. However, adapting models to represent quantitative concepts, such as object counts, poses unique challenges as it involves precise numeric representation rather than qualitative identity. In contrast, our work focuses on achieving precise, domain-generalizable object quantification across varied styles and contexts without retraining, extending personalized image generation to handle quantitative control across different visual styles and settings.
Optimizing Prompt Correspondence. Recent studies have improved the alignment between text prompts and generated images. Prompt-to-Prompt [17] manipulates textual encoder attention for object replacement, while Null-inversion [36] allows similar edits on real images. Attend-and-Excite [5] ensures all objects in a prompt are depicted by optimizing attention maps. However, these methods generally do not address numeric accuracy in object counts, as prompt manipulation alone often fails to ensure precise quantities in diffusion outputs. Our work addresses this gap by introducing meta-optimized prompts that enable reliable control over object quantities across varied styles and contexts, establishing a foundation for robust object quantification in diverse visual settings.
@@
Domain Generalization. Domain generalization (DG) aims to learn knowledge that transfers well to unseen target domains, allowing models to generalize effectively without access to target domain data during training. DG is typically studied under two main settings: multi-source domain generalization and single domain generalization (SDG). Multi-source DG utilizes multiple source domains to enhance robustness, while SDG—considered more realistic yet challenging—relies on a single source domain, making it harder due to limited data diversity. Early approaches to DG drew inspiration from domain adaptation methods, focusing on domain alignment to learn domain-invariant features. Techniques such as adversarial training [25, 53] and distribution distance minimization [32, 33] were employed to reduce feature bias across multiple source domains. Additionally, self-supervised learning [4, 26] and meta-learning [9, 55] have been explored to improve generalization in DG tasks. However, to the best of our knowledge, no prior work has specifically addressed domain-generalizable object quantification in text-to-image diffusion models. Our work introduces a novel setting and methodology to achieve domain-invariant object quantification across diverse and unseen domains. This new setting expands the scope of domain generalization research to the area of text-to-image generation, where accurate object quantification across varied domains is crucial.
Meta-Learning. Meta-learning, or learning to learn, seeks to enable rapid and efficient adaptation to new tasks by leveraging prior learning experiences, as reviewed by Hospedales et al. [21]. It has been applied across diverse areas, including designing adaptive loss functions [54, 2, 3], creating task-specific model initializations [12], and advancing few-shot learning techniques [28, 58, 56, 9, 10]. Meta-learning approaches are generally categorized into metric-based [56, 58, 61, 29], memory-based [35, 52, 37, 38, 20, 11], and gradient-based methods [48, 16, 39, 31]. Since the introduction of Model-Agnostic Meta-Learning (MAML) [12], gradient-based approaches have gained popularity, though they often encounter challenges with meta-overfitting due to limited meta-training tasks [1, 64, 69, 22, 23, 27]. Overall, meta-learning serves as a crucial foundation for our approach, enabling effective domain generalization in object quantification by facilitating adaptive prompt optimization across diverse text-to-image generation tasks.
@@ 3 Preliminaries
Text-Conditioned Diffusion Process. The diffusion process generates an image $x\sim p(x)$ from random noise through an iterative denoising procedure. Given an image $x_{0}\sim p(x)$, training involves creating noisy samples $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon_{t}$, where $\alpha_{t}$ are hyperparameters and $\epsilon_{t}\sim\mathcal{N}(0,\text{I})$. A neural network is trained to predict the added noise $\epsilon_{t}$ using the following objective:
$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t)-\epsilon_{t}\|_{2}^{2}\right].$ (1)
For conditional denoising, a similar objective is applied, with each step conditioned on an external factor $y$:
$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t,y)-\epsilon_{t}\|_{2}^{2}\right].$ (2)
In text-to-image diffusion, training involves reconstructing images conditioned on textual descriptions, which guide the denoising process to match the text. This requires a large dataset to generalize mappings between text and object quantities accurately. Some approaches modify noise predictions to align with specific object counts [66], although these often require counting models capable of handling noisy, intermediate data. In our work, we propose a method that adjusts the text conditioning embeddings using feedback from an off-the-shelf counting model.
Meta-Learning. Meta-learning aims to build models capable of quickly adapting to new tasks with limited data by utilizing prior experiences across various tasks [12]. Let $\mathcal{D}$ represent a meta-training dataset composed of training and validation sets across a range of tasks $\mathcal{T}$, i.e., $\mathcal{D}=\{\{D_{i}^{\text{tr}},D_{i}^{\text{val}}\}\}_{i\in\mathcal{T}}$, where $D_{i}^{\text{tr}}$ and $D_{i}^{\text{val}}$ are the training and validation datasets for the $i$-th task. Meta-learning can be formulated as a bi-level optimization problem:
$\min_{\phi}$ $\;\sum_{i\in\mathcal{T}}\mathcal{L}_{\text{val}}({\hat{\theta_{i}}}(\phi);D^{\text{val}}_{i})$ (3)
s.t. $\;{\hat{\theta_{i}}}(\phi)=\operatorname*{arg\,min}_{\theta_{i}}\mathcal{L}_{\text{train}}(\theta_{i};\phi,D^{\text{tr}}_{i}),\forall i\in\mathcal{T},$ (4)
where $\mathcal{L}_{\text{val}}$ and $\mathcal{L}_{\text{train}}$ represent the validation and training losses for the upper- and lower-level optimizations, respectively. In this setup, $\theta_{i}$ denotes the task-specific parameters for the $i$-th task, while $\phi$ are the meta-parameters shared across tasks. The lower-level optimization (Eq. (4)) adapts task-specific parameters $\theta_{i}$ for each task using the meta-parameters $\phi$ and the task-specific training set $D_{i}^{\text{tr}}$. Meanwhile, the upper-level optimization (Eq. (3)) aims to find meta-parameters $\phi$ that minimize the aggregated validation loss across tasks based on the optimized task-specific parameters $\hat{\theta_{i}}(\phi)$. Our work leverages this meta-learning framework to achieve domain-generalizable object quantification in text-to-image diffusion models, employing a dual-loop optimization approach to refine style-invariant prompts across diverse domains.
@@ 4 Methods
Figure 2: QUOTA’s dual-loop meta-optimization. The inner loop optimizes prompt parameters $e_{\text{count}}$ and $e_{\text{style}}$ across meta-train domains to adapt to domain-specific prompts, with losses computed using YOLOv9 for quantification accuracy and CLIP for semantic alignment. The outer loop refines these parameters on a meta-test domain, enhancing the model’s generalization to unseen domains. This dual-loop structure enables consistent object quantification and domain adaptation across diverse domains without the need for retraining.
In this section, we introduce quantifying objects with text-to-image models for any domain (QUOTA), our approach for quantifying object quantities across diverse domains in text-to-image generative models. QUOTA leverages a dual-loop meta-learning framework to create domain-invariant prompts that guide object quantification without retraining, making it adaptable to new and unseen domains. Through prompt learning, QUOTA optimizes domain-specific and quantity-specific parameters to accurately adjust object quantities based on textual prompts, even when encountering new stylistic variations. Our framework operates on images generated by existing text-to-image models, such as SDXL [41], using object quantities inferred by a pre-trained detection model as feedback for optimization.
@@ 4.1 Domain-generalizable object quantification
Accurately controlling the quantity of objects in text-to-image generative models is essential for producing realistic images, especially when adapting to diverse domains. To address the challenge of domain generalization, we introduce a novel training framework where multiple domains are leveraged to optimize the text-to-image generation process for consistent object quantification. Let $\mathcal{D}=\{\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}\}$ represent the set of domains available in the training phase, where $\mathcal{D}_{\text{train}}$ denotes the meta-train domains and $\mathcal{D}_{\text{test}}$ represents a held-out meta-test domain. During each iteration, we randomly select one domain from the training domains as the meta-test domain $\mathcal{D}_{\text{test}}$, while the remaining domains are assigned as the meta-train domains $\mathcal{D}_{\text{train}}$. This randomized selection allows the model to experience different combinations of training and testing domains, enhancing its robustness and adaptability to unseen domains. During training, images are generated based on text prompts in each domain using a pre-trained diffusion model, such as SDXL [41], without access to ground truth images. This setting enables us to simulate diverse training scenarios, preparing the model for domain shifts.
In the testing phase, our goal is to generalize to an unseen domain $\mathcal{D}_{\text{unseen}}$, where the model is expected to quantify objects accurately in images generated by the text-to-image model. To evaluate the object counts in these generated images, we employ two pre-trained models: YOLOv9 [62] and CLIP-S [44]. YOLOv9 serves as an object detector, providing a quantitative measure of the number of objects present in the image. Meanwhile, CLIP-S offers an alignment score between the generated image and the text prompt, helping to assess how well the generated image adheres to the prompt’s specifications, including object count. Notably, since the diffusion process does not inherently support direct control over object counts, we rely on a pre-trained object counting model that operates on completed images rather than intervening in the intermediate denoising steps. Thus, our method focuses on refining the domain and quantification parameters of the prompt, allowing it to adapt to the generated image characteristics across various domains and ensure accurate quantify object in the final output.

### v2

@@ Abstract
We tackle the problem of quantifying the number of objects by a generative text-to-image model. Rather than retraining such a model for each new image domain of interest, which leads to high computational costs and limited scalability, we are the first to consider this problem from a domain-agnostic perspective. We propose QUOTA, an optimization framework for text-to-image models that enables effective object quantification across unseen domains without retraining. It leverages a dual-loop meta-learning strategy to optimize a domain-invariant prompt. Further, by integrating prompt learning with learnable counting and domain tokens, our method captures stylistic variations and maintains accuracy, even for object classes not encountered during training. For evaluation, we adopt a new benchmark QUANT-Bench specifically designed for object quantification in domain generalization, enabling rigorous assessment of object quantification accuracy and adaptability across unseen domains in text-to-image generation. Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new benchmark for efficient and scalable text-to-image generation for any domain.
@@ 1 Introduction
Text-to-image generative models [37, 22, 31, 4] have achieved remarkable success in generating high-quality images from textual descriptions. However, accurately controlling the count of objects in these generated images remains challenging. Recently, Zafar et al. [35] introduced prompt learning to refine the object counting prompts. While effective in controlled scenarios, this method struggles to maintain accurate object counts when confronted with varying domains. Similarly, CountGen [1] improves counting accuracy by guiding the denoising process through internal object representations; however, it is restricted to a single domain and depends on additional modules for layout prediction. As a result, achieving domain-generalizable object quantification remains a crucial and unresolved challenge across various generative models, including diffusion and flow-matching models. Figure 1 illustrates this problem, showing how standard generative models like SDXL [23] struggle with accurate object quantification across different domains, whereas our approach provides consistent and accurate results for any domain.
Figure 1: Quantifying objects across domains. In each column, one domain is held out as the unseen test domain (shown in the column title), while the remaining three domains are used for training. For example, in the first column Photo is the test domain and the models are trained on Sketch, Cartoon, and Painting. Red boxes indicate incorrect object counts. SDXL [23] and IoCo [35] struggle with consistent quantification across domains. In contrast, QUOTA accurately controls object counts across all domains, including unseen ones.
@@
Domain generalization in machine learning aims to create models that perform well on unseen domains, reducing the need for retraining and enabling greater adaptability across diverse settings [38]. A wide range of approaches has been developed for this purpose, including methods that focus on learning domain-invariant features and leveraging diverse training environments [16, 5, 34]. Among these approaches, meta-learning has emerged as a promising technique, enabling models to adapt to new tasks by utilizing prior knowledge from related tasks [13, 32, 6]. Despite these advancements, there has been no exploration of meta-learning for domain-generalizable object quantification in text-to-image generation. In this paper, we fill this gap by proposing a novel meta-learning framework designed to enable robust object quantification across diverse and unseen domains within text-to-image models.
@@
Specifically, we make the following three contributions in this paper. First, we introduce a domain-agnostic setting for object quantification in generative text-to-image models, emphasizing the need for precise, domain-invariant object quantification. This setting is essential for adapting models to diverse and unseen prompts without retraining, addressing both scalability and efficiency challenges. Second, we propose a dual-loop meta-learning strategy for prompt optimization, designed to make prompts domain-invariant. In the training phase, we divide text prompts into meta-train and meta-validation domains. In the inner loop, we optimize prompt parameters on the meta-train domain to meet specific domain requirements. The outer loop refines these parameters on the meta-validation domain, allowing the model to learn a prompt that generalizes across unseen domains. This design mimics real-world scenarios where models must handle diverse domain shifts without prior exposure, ensuring adaptability to novel distributions in text-to-image generation. Third, we incorporate prompt learning with learnable quantification and domain tokens, enabling the model to capture stylistic variations effectively while preserving object quantification accuracy across domains. This approach allows for robust adaptation to new domains and unseen object classes without modifying the model architecture or retraining.
@@
For evaluation, we adopt a new benchmark QUANT-Bench specifically designed for object quantification in domain generalization, enabling rigorous assessment of accuracy and adaptability across unseen domains in text-to-image generation. Comparative evaluations against state-of-the-art methods demonstrate that QUOTA outperforms existing models in both object quantification accuracy and semantic alignment, setting a new benchmark for domain-generalizable text-to-image generation, enabling accurate object quantification across any domain without retraining.
@@ 2 Related Work
Text-to-image generation. The field of image generation has experienced significant evolution, transitioning from initial Generative Adversarial Network (GAN)-based approaches [8] to the more recent diffusion-based models [12]. Early GAN methodologies laid the groundwork for generating images from textual descriptions, but often faced challenges related to training stability and output quality [8]. In contrast, diffusion models have emerged as a robust alternative, demonstrating enhanced performance in producing high-fidelity images conditioned on text prompts [12]. Notably, models like DALL·E 2 [24] and Imagen [27] have showcased the capability to generate photorealistic images from textual inputs. Concurrently, flow matching models have been introduced, offering advantages in training stability and sample quality by aligning continuous normalizing flows with target distributions [19]. Our approach builds upon these advancements in text-conditioned image generation, focusing on enhancing object counting consistency across diverse domains.
Personalized image generation. Personalized image generation aims to fine-tune models for specific concepts using limited examples. Textual Inversion [7] learns a new token with just a few images, while DreamBooth [26] fine-tunes diffusion models for personalized content. However, adapting models for quantitative concepts, such as object counting, remains challenging as it requires precise numerical control rather than qualitative representation. Recent works have improved prompt-image alignment through techniques like Prompt-to-Prompt [9] for object manipulation, Null-inversion [20] for real-image edits, and Attend-and-Excite [3] to ensure prompt consistency. Our work introduces meta-optimized prompts that enable accurate, domain-generalizable object quantification without retraining, ensuring reliable control over object counts across diverse styles and contexts for consistent text-to-image generation.
Object counting benchmarks. Several benchmarks have been proposed for evaluating object counting in text-to-image generation, including DrawBench[27], CoCoCount[1], and YOLO-Count [36]. DrawBench evaluates multiple aspects of image generation, including cardinality, but does not address domain generalization. CoCoCount focuses on counting accuracy within specific datasets (e.g., COCO), but remains domain-specific and does not generalize across unseen domains. On the other hand, YOLO-Count introduces benchmarks like OpenImg7-New and Obj365-New, which assess a model’s ability to generalize to unseen categories. In contrast, our QUOTA benchmark focuses on precisely generating objects across unseen domains without retraining, providing a more scalable solution for real-world applications.
@@
Domain generalization. Domain generalization (DG) aims to enable models to generalize to unseen domains without access to target domain data during training. It is typically studied under two settings: multi-source DG, which leverages multiple source domains for robustness, and single-domain DG (SDG), which is more realistic but challenging due to limited data diversity. Early DG approaches focused on domain-invariant feature learning via adversarial training [14, 29], distribution alignment [17, 18], and self-supervised learning [2, 15]. Meta-learning [5, 30] has also been explored to enhance generalization. However, no prior work has specifically addressed domain-generalizable object quantification in text-to-image diffusion models. We introduce a novel setting and methodology to achieve domain-invariant object quantification across diverse and unseen domains.
@@ 3 Preliminaries
Iterative object count optimization (IoCo). IoCo by Zafar et al. [35] introduces an iterative method to optimize object count accuracy in text-to-image models. Given a generated image $x$, an object category $c$, and a target object count $N$, their framework refines the generated output by minimizing the following objective:
$\min_{\theta}\mathcal{L}_{\text{counting}}(x,c,N)+\lambda\mathcal{L}_{\text{semantic}}(x,c,N),$ (1)
where the counting loss ensures the estimated count aligns with the target:
$\mathcal{L}_{\text{counting}}(x,c,N)=\left|g_{\text{cnt}}(x,c)-N\right|,$ (2)
with $g_{\text{cnt}}(x,c)$ representing the estimated number of objects in the generated image. To maintain the semantic integrity of the image, a CLIP-based regularization is used:
$\mathcal{L}_{\text{semantic}}(x,c,N)=\text{CLIP}(x,\text{``A photo of }N\text{ }c\text{''}).$ (3)
IoCo iteratively updates a learnable counting token $e_{\text{cnt}}$ to adjust the model’s internal representation of quantity:
$e_{\text{cnt}}\leftarrow e_{\text{cnt}}+\alpha\frac{\nabla_{e_{\text{cnt}}}\mathcal{L}(x,c,N)}{\|\mathcal{L}(x,c,N)\|_{2}}.$ (4)
Meta-learning for domain generalization. In this work, meta-learning aims to develop models capable of rapidly adapting to new tasks with limited data by leveraging knowledge from multiple tasks [6]. In the context of domain generalization, meta-learning seeks to train models that generalize well to unseen domains by simulating domain shifts during training [16]. Let $\mathcal{S}$ represent a training dataset comprising training and validation sets across various source domains $\mathcal{S}$, i.e., $\mathcal{S}=\{\mathcal{S}^{\text{tr}},\mathcal{S}^{\text{val}}\}$, where $\mathcal{S}^{\text{tr}}$ and $\mathcal{S}^{\text{val}}$ are disjoint subsets of the source domain set $\mathcal{S}$, with the former used for training and the latter for validation. The meta-learning process for domain generalization can be formulated as a bi-level optimization problem:
$\min_{\phi}$ $\;\sum_{s^{\text{val}}\in\mathcal{S}^{\text{val}}}\mathcal{L}_{\text{val}}({\theta_{s^{\text{val}}}^{*}};s^{\text{val}}).$ (5)
s.t. $\;{\theta_{s^{\text{val}}}^{*}}=\operatorname*{arg\,min}_{\theta_{s^{\text{tr}}}}\mathcal{L}_{\text{train}}({\theta_{s^{\text{tr}}}};s^{\text{tr}}),\quad s^{\text{tr}}\in\mathcal{S}^{\text{tr}},$ (6)
where $\mathcal{L}_{\text{val}}$ and $\mathcal{L}_{\text{train}}$ denote the validation and training losses at the upper and lower levels, respectively. Here, $\theta_{s^{\text{tr}}}$} represents the domain-specific parameters for source domain $s^{\text{tr}}$, and $\phi$ denotes the shared meta-parameters. The lower-level optimization (Eq. (6)) adapts $\theta$ to each source domain using the training set $\mathcal{S}^{\text{tr}}$ and the meta-parameters $\phi$. The upper-level optimization (Eq. (5)) updates $\phi$ by minimizing the combined validation loss across source domains for the optimized domain-specific parameters $\theta_{s^{\text{val}}}^{*}$}. During testing, the trained model is evaluated on data from unseen target domains $\mathcal{T}$, which were not accessible during training. The goal is to assess the model’s ability to generalize to these novel domains without additional fine-tuning.
@@ 4 Methods
In this section, we introduce quantifying objects with text-to-image models for any domain (QUOTA). QUOTA leverages a dual-loop meta-learning framework to create domain-invariant prompts that guide object quantification without retraining, making it adaptable to new and unseen domains. Through prompt learning, QUOTA optimizes domain-specific and quantity-specific parameters to accurately adjust object quantities based on textual prompts, even when encountering new stylistic variations. Our framework operates on images generated by existing text-to-image models, such as SDXL [23], using object quantities inferred by a pre-trained detection model as feedback for optimization.
@@ 4.1 Domain-generalizable object quantification
Accurately quantifying objects in text-to-image generative models is essential for ensuring alignment with specified numerical constraints. However, achieving consistent object counts across diverse domains is challenging due to the inherent variability in domain-specific styles and the lack of explicit numerical control in diffusion-based generation.
Let $\mathcal{S}^{\text{tr}}$ represent a set of source domains used during training. Existing methods for object quantification in text-to-image diffusion models primarily focus on optimizing count accuracy within a fixed domain $s^{\text{tr}}\in\mathcal{S}^{\text{tr}}$ but overlook the challenge of generalization. These approaches assume that the model is trained and evaluated within the same domain, failing to account for domain shifts encountered when generating images in unseen distributions. As a result, a model trained on $\mathcal{S}^{\text{tr}}$ may overfit to domain-specific characteristics, leading to inconsistent quantification performance when applied to novel styles.
In this work, we formally introduce the problem of domain-generalizable object quantification in text-to-image diffusion models. Our goal is to ensure that object quantification remains reliable when applied to unseen target domains $\mathcal{T}$, where object distributions, styles, and scene compositions may differ considerably from $\mathcal{S}^{\text{tr}}$. To the best of our knowledge, this is the first study that investigates how to optimize object counting mechanisms in a way that remains robust across unseen domains. Without explicitly addressing domain generalization, object quantification methods risk becoming unreliable when applied to new domains, limiting their practical utility in real-world scenarios.
Figure 2: QUOTA’s dual-loop meta-optimization. The inner loop optimizes prompt parameters $e_{\text{count}}$ and $e_{\text{style}}$ across meta-train domains (Cartoon and Sketch) to adapt to domain-specific prompts, with losses computed using YOLOv9 for quantification accuracy and CLIP for semantic alignment. The outer loop refines these parameters on a meta-validation domain (Painting), enhancing the model’s generalization to unseen domains (Photo). This dual-loop structure enables consistent object quantification and domain adaptation across diverse domains without retraining.